Foxx Corporation acquired all of Greenburg Company’s outstanding stock on January 1, 2016, for $771,000 cash. Greenburg’s accounting records showed net assets on that date of $542,000, although equipment with a 10-year life was undervalued on the records by $168,000. Any recognized goodwill is considered to have an indefinite life.

Greenburg reports net income in 2016 of $112,000 and $135,000 in 2017. The subsidiary declared dividends of $20,000 in each of these two years.

Account balances for the year ending December 31, 2018, follow. Credit balances are indicated by parentheses.

Foxx Greenburg
Revenues $ (1,108,000 ) $ (724,000 )
Cost of goods sold 138,500 181,000
Depreciation expense 410,000 381,000
Investment income (20,000 ) 0
Net income $ (579,500 ) $ (162,000 )
Retained earnings, 1/1/18 $ (1,224,000 ) $ (449,000 )
Net income (579,500 ) (162,000 )
Dividends declared 120,000 20,000
Retained earnings, 12/31/18 $ (1,683,500 ) $ (591,000 )
Current assets $ 365,000 $ 128,000
Investment in subsidiary 771,000 0
Equipment (net) 1,074,000 766,000
Buildings (net) 954,000 562,000
Land 704,000 131,000
Total assets $ 3,868,000 $ 1,587,000
Liabilities $ (1,284,500 ) $ (696,000 )
Common stock (900,000 ) (300,000 )
Retained earnings (1,683,500 ) (591,000 )
Total liabilities and equity $ (3,868,000 ) $ (1,587,000 )
  1. Determine the December 31, 2018, consolidated balance for each of the following accounts:

Depreciation Expense Buildings
Dividends Declared Goodwill
Revenues Common Stock
Equipment
  1. How does the parent's choice of an accounting method for its investment affect the balances computed in requirement (a)?

  2. Which method of accounting for this subsidiary is the parent actually using for internal reporting purposes?

  3. Determine parent's investment income for 2018 under partial equity method and equity method.

  4. What would be Foxx’s balance for retained earnings as of January 1, 2018, if each of the following methods had been in use?

  • Initial value method.
  • Partial equity method.
  • Equity method.

rev: 11_02_2017_QC_CS-107892, 09_28_2019_QC_CS-183150, 02_04_2020_QC_CS-198743
Answer #1

Dear Student,

If any doubts then feel free to ask

Part A

Purchase price

771000

Book value (given)

(542000)

Price in excess of book value

229000

Life

Annual Excess amortizations

Allocation to equipment based on difference in market value and book value

168000

10

16800

Goodwill

61000

indefinite

Total

$229000

$16800

Consolidated balances

Depreciation expense

$807800

Dividends declared

120000

Revenues

1832000

Equipment

1957600

Buildings

1516000

Goodwill

61000

Common stock

900000

Depreciation expense = book value of Foxx + book value of Greenburg+amortization expense = 410000+381000+16800=$807800

Dividends declared = book value of Foxx = 120000

Revenues = book value of Foxx + book value of Greenburg = 1108000+724000=$1832000

Equipment= book value of Foxx+ book value of Greenburg + excess allocation+ amortization expense(3 years) = 1074000+766000+168000-(16800*3) = 1957600

Buildings= book value of Foxx + book value of Greenburg = 954000+562000=$1516000

Goodwill= purchase price of net assets-book value of net assets - excess allocated to equipment = 771000-542000-168000=$61000

Common stock = book value of Foxx = 900000

Part B

Option B

No, doesn't affect consolidated totals but only internal reporting of parent.

The parent's choice for accounting methods for its investment does not affect the consolidated balances of depreciation, dividends declared, equipment, revenue, common stock, goodwill, buildings, revenues.

Part C

Option A

Initial value method

Currently, the parent company is currently using the initial value method for reporting. The parent's investment investment in subsidiary is equal to $771000 which is its original cost. Moreover investment income is equal to dividends declared by subsidiary.

Part D

Investment income

Partial equity method

$162000

Equity method

$145200

Partial equity method = net income of Greenburg= 162000

Equity method = net operating income-depreciation = 162000-16800 = 145200

Part E

Retained earnings

Initial value method

$1224000

Partial equity method

$1431000

Equity method

$1397400

Initial value method = Beginning retained earnings of Foxx = $1224000

Partial equity method = Beginning retained earnings+ net income 2016 + net income 2017 - dividends = 1224000+112000+135000-(20000*2) = $1250000

Equity method = Beginning retained earnings+ net income 2016 + net income 2017 - dividends - depreciation expense 2016 - depreciation expense 2017 = 1224000+112000+135000-(20000*2)-16800-16800 = 1397400